A. A life insurance company will sell a Php250,000 one-year term life insurance policy for members of armed forces of the Philippines for a premium of Php500.00. Let X denote the gain from the insurance company. Based on the collected data of the company, a member of the armed forces police has a 99.96% chance of surviving within one year. Let X the net gain of an insurance company from a single policy holder.



1. Complete the table below showing the probability distribution of X.



X 500 -249,500


P(x)



2. Find the expectation If you play a bet and Interpret the result.



E(X)=



3. Interpretation:​

1
Expert's answer
2022-04-01T05:02:35-0400

"EX=500\cdot 0.9996-249500\cdot 0.0004=400"

The expectation is positive because company wants to gain money.